Windvi 0.65 Features

Fabrice Popineau 21/07/1998

Contents

1 Introduction 2 2 Postscript inclusions 3 3 Arbitrary Postscript code 4 4 TPIC specials 5 5 Transformations 6 6 The world of color 7 7 The XY-Pic package 8 8 The PSTricks package 9 9 Is color correctly handled ? 11 10 Background color for the whole page 13 11 External commands and inclusions 15

1 Introduction

Many of these examples are taken from the LaTeX Graphics Companion. First, we check the color text behavior:

green text red text yellow text magenta text cyan text

This is the default text.

2 Postscript inclusions

Various effects:

left  right left  right left  right

The same file, but in a rotated box:

left right

You can include the compressed versions too:

left  right left  right

The first one is GZip’ed, the second one is BZip2’ed.

3 Arbitrary Postscript code

The following figure, Fig. 1, is an example of raw Postscript being sent to the driver. It has been taken from the dvips manual.

Figure 1: Postscript code directly from a special command.

This code lead to the previous figure:

  \vspace{2in}
  \vbox to 100bp{
    \special{" newpath 000 000 moveto 100 100 lineto 394 0 lineto
      closepath gsave 0.8 setgray fill grestore stroke}\vfil}

4 TPIC specials

A TPiC trial:

C B B A  A C A C B A C B Start  1  2  3  4  *

End of TPic test.

And the pspicture environment:

5 Transformations

Here the text should be rotated, but given this is text, and that this material is not processed by ghostscript, the text is not rotated.

However, under Windows NT, there is an opportunity to render this and this is done now. So only NT users will see the actual text.

A  A  A  A  A

TUG96 in Russia TUG96 in Russia TUG96 in Russia TUG96 in Russia TUG96 in Russia And with tables :

1 2 3
a b c

6 The world of color

1. magenta cmyk black

2. predefined blue gray text

Black text, blue background, red frame White text, blue background, red frame Green text, blue background, red frame

Fun with color Fun with color

Fun with color Fun with color

7 The XY-Pic package

[figure]

8 The PSTricks package

RAMSU

SRawData Scombine ⊕

⊕ SGenerateData

RModes RNormal RVertical RLatched RTripped ROther This is the same tree, but rotated. Only NT users will see the glyphs at the right place. Win9x will see the Postscript code at the right place, but the glyphs displaced:

RAMSU

SRawData Scombine ⊕

⊕ SGenerateData

RModes RNormal RVertical RLatched RTripped ROther

9 Is color correctly handled ?

Something that is higly desirable : being able to jump to any page, and be placed in the correct color stack state. This is possible because Windvi does pre-scanning of all specials.

What will happen if we break some page across color text ? Will we get the expected result ?

10 Background color for the whole page

Test of the background for the windvi program.

Is the background restored to white ? I hope not ! Because the \pagecolor{} command is sticky through out

the document.

11 External commands and inclusions

This is an inclusion of a .jpg image thanks to jpeg2ps.exe.

Figure 2: The author.

Oh ! By the way : that’s me. This was the easiest jpeg picture to find. Now trying .bmp files:

Figure 3: Some bmp file.

And Windows Enhanced Metafiles: Figure 4: Some emf file.